# Spectral Graph Pruning Against Over-Squashing and Over-Smoothing

### v1

@@ 2 Theoretical Insights Into Spectral Rewiring
Note that $\widetilde{\mathcal{G}^{+}}$ decreases the spectral gap, while $\mathcal{G}^{+}$ increases it relative to $\mathcal{G}$. However, the label configuration of $\widetilde{\mathcal{G}^{+}}$ seems more advantageous because for the changed nodes, the number of neighbors of the same class label remain in the majority in contrast to $\mathcal{G}^{+}$. Still, the spectral gap increase seems to aid the learning task compared to the spectral gap decrease.
@@ 3 Braess-inspired graph rewiring
In the following, we propose two algorithmic approaches to perform spectral rewiring. Our main proposal is computationally more efficient and more effective in spectral gap approximation than baselines, as we also showcase in Table 1. The other approach based on Eldan’s Lemma is also analyzed, as it provides theoretical guarantees for edge deletions. However, it does not scale well to larger graphs.
@@ 3.1 Graph Rewiring with Proxy Spectral Gap Updates
Time Complexity of ProxyAdd. The run time analysis consists of the same elements as the edge deletion algorithm. The key distinction is that the ranking is conducted on the complement of the graph’s edges, $\bar{\mathcal{E}}$. Since the set of missing edges is usually larger than the existing edges in real world settings, to save computational overhead for large graphs it is possible to only sample a constant amount of them.
@@ 3.2 Graph Rewiring with Eldan’s Criterion
Lemma 2.1 states a sufficient condition for deleting edges for the Braess phenomenon to occur on a graph’s spectral gap. It naturally defines a scoring function of edges to rank them according to their potential to maximize the spectral gap based on the function $g$. However, the computation of this ranking is significantly more expensive that other considered algorithms, as each scoring operation needs access to the leading eigenvector of the perturbed graph with an added or deleted edge. In case of edge deletions, we also need to approximate the spectral gap similar to our Proxy algorithms. Furthermore, the involved projection $\mathcal{P}_{f}$ is a dot product of eigenvectors and therefore runs in $\mathcal{O}(|\mathcal{V}|)$. Even though this algorithm does not scale well to large graphs without focusing on a small random subset of candidate edges, we still consider it as baseline, as it defines a more conservative stopping criterion to assess when we should stop deleting edges. The precise algorithms are stated in Appendix C.
@@ 3.3 Approximation Quality
Runtime analysis. Furthermore, Table 1 provides evidence for the fact that the proposed algorithms ProxyAdd and ProxyDelete can be computationally faster than baselines. See Appendix F for a detailed discussion on runtimes.
Table 1: Runtimes for 50 edge modifications in seconds.
Method Cora Citeseer Chameleon Squirrel
FoSR 4.69 5.33 5.04 19.48
SDRF 19.63 173.92 17.93 155.95
ProxyAdd 4.30 3.13 1.15 9.12
ProxyDelete 1.18 0.86 1.46 7.26
@@ 4 Experiments
4.1 Long Range Graph Benchmark
The Long Range Graph Benchmark (LRGB) was introduced by (Dwivedi et al., 2023). The graph datasets have been created specifically to test if message passing graph neural networks can capture long range interactions between nodes. Thus, these datasets serve as a nice benchmark to test methods that claim to mitigate oversquashing. We compare our proposed ProxyAdd and ProxyDelete methods with DRew (Gutteridge et al., 2023), a recently proposed strong baseline for addressing oversquashing with GCN as our backbone architecture in Table 2. We adopt the experimental setting of Tönshoff et al. (2023). The values for DRew are taken from the paper.
We test on the following datasets and tasks: 1) PascalVOC-SP - Semantic image segmentation as a node classification task operating on superpixel graphs. The metric choosen to evalauate is F1. 2) Peptides-func - Peptides modeled as molecular graphs. The task is graph classification. The metric is AP. 3) Peptides-struct - Peptides modeled as molecular graphs. The task is to predict various molecular properties, hence a graph regression task. The metric is MAE.
@@ 4.1 Long Range Graph Benchmark
The top performance is highlighted in bold. Evidently, our proposed rewiring methods outperform DRew (Gutteridge et al., 2023) on PascalVOC and Peptides-struct and comparably on Peptides-func datasets. In Table 10 we also provide node classification results for other commonly used datasets comparing with different rewiring strategies. For a fair comparison, we also perform graph classification for adding edges, since FoSR (Karhadkar et al., 2023) was primarily tested on this task. Table 11 highlights the utility of our methods on all considered datasets.
Method Cora Citeseer Pubmed
Metrics GS WS Acc GS WS Acc GS WS Acc
@@
Platonov et al. (2023) point out that most progress on heterophilic datasets is unreliable since many of the used datasets have drawbacks, including duplicate nodes in Chameleon and Squirrel datasets, which lead to train-test data leakage. The sizes of the small graph sizes also lead to high variance in the obtained accuracies. Consequently, we also test our proposed algorithms on 3/5 of their newly introduced larger datasets and use GCN (Kipf & Welling, 2017) and GAT (Veličković et al., 2018) as our backbone architectures. Furthermore, we also increase the number of layers and check how our method fares at higher depths that potentially increases over-smoothing. To that end, we adopt the code base and experimental setup of Platonov et al. (2023); the datasets are divided into 50/25/25 split for train/test/validation respectively. The test accuracy is reported as averaged over 10 runs. To facilitate training deeper models, skip connections and layer normalization is also used. We compare FoSR (Karhadkar et al., 2023) and our proposals based on the Eldan criterion as well as ProxyAdd and ProxyDelete in Tables 4,5,6. The top peformance is highlighted in bold. Evidently, for increasing depth, even though the GNN performance should degrade because of over-smoothing, we achieve a significant boost on accuracy compared to baselines, which we attribute to the fact that our methods delete inter-class edges —thus slowing down detrimental smoothing. The empirical runtimes of different rewiring algorithms are compared in Table 14.
@@
Table 5: Node Classification on Amazon-Ratings dataset.
@@
Table 6: Node Classification on Minesweeper dataset.
@@
In Sections §2.2 and 4, we have shown that graph pruning can improve generalization, mitigate over-squashing and also help slow down the rate of smoothing. Can we also use our insights to find lottery tickets (Frankle & Carbin, 2019)?
@@
To what degree is graph pruning feature data dependent? The first extension of the Lottery Ticket Hypothesis to GNNs, called Unified Graph Sparsification (UGS) Chen et al. (2021), prunes connections in the adjacency matrix and model weights that are deemed less important for a prediction task. Note that UGS relies on information that is obtained in computationally intensive prune-train cycles that takes into account the data and the associated masks. In the context of GNNs, the input graph plays a central role in determining a model’s performance at a downstream task. Naively pruning the adjacency matrix without characterizing what constitutes important edges is a pitfall we would want to avoid (Hui et al., 2023), yet resorting to expensive train-prune-rewind cycles to identify importance is also undesirable. This brings forth the questions: To what extent need the pruning criteria depend on the data under consideration? Is it possible to formulate a data/feature agnostic pruning criterion that optimizes a more general underlying principle to find lottery tickets? Morcos et al. (2019) and Chen et al. (2020) show, in the context of computer vision and natural language processing respectively, that lottery tickets can have universal properties that can even provably (Burkholz et al., 2022) transfer to related tasks. However, this universality is only achieved at moderate sparsity levels. In the high sparsity regime, architectures seem to require some form of task specialization.
@@
Lottery tickets that rely on the spectral gap. However, even specialized structures need to maintain and promote information flow through its connections. This fact has inspired works like Pal et al. (2022); Hoang et al. (2023) to analyze how well lottery ticket pruning algorithms maintain the Ramanujan graph property of bipartite graphs, which is intrinsically related to the Cheeger constant and thus the spectral gap. They have further shown that rejecting pruning steps that would destroy a proxy of this property can positively impact the training process. In the context of GNNs, we show that we can base the graph pruning decision even entirely on the spectral gap, but rely on a computationally cheaper approach to obtain a proxy. By replacing the magnitude pruning criterion for the graph with the Eldan criterion and ProxyDelete to prune edges, in principle, we can avoid the need for additional data features and labels. This has the advantage that we could also prune the graph at initialization and thus benefit from the computational savings from the start. We use our proposed methods to prune the graph at initialization to requisite sparsity level and then feed it to the GNN where the weights are pruned in an iterative manner. Our results are presented in Table 17, where we compare IMP based UGS (Chen et al., 2021) with our methods for different graph and weight sparsity levels. Note that, even though our method does not take into account any feature information into account and prunes purely based on the graph structure, our results are comparable, and for datasets like Pubmed we even slightly outperform the baseline. Table 3 shows results for jointly pruning the graph and parameter weights, which leads to better results due to potential positive effects of overparameterization on training (Gadhikar & Burkholz, 2024).
@@
Stopping criterion. The advantage of using spectral gap based pruning (especially Eldan criterion) is patent: It helps identify problematic edges that cause information bottlenecks and provides a framework to prune those edges. Unlike UGS, our proposed framework also has the advantage that we can couple the overall pruning scheme with a stopping criterion that follows naturally from our set-up. We stop pruning the input graph when no edges that satisfy our criterion are available anymore.
@@ 5 Conclusion
Our work connects two seemingly distinct branches of the literature on GNNs: rewiring graphs to mitigate over-squashing and pruning graphs for lottery tickets to save computational resources. On the one hand, we highlight that, contrary to the standard rewiring practice, not only adding but also pruning edges can increase the spectral gap of a graph. We exemplify theoretically and experimentally that this fact implies that we can address over-squashing and over-smoothing simultaneously. On the other hand, these results explains how pruning graphs moderately can improve the generalization performance of GNNs, in particular for heterophilic learning tasks. To utilize these insights, we have proposed a computationally efficient graph rewiring framework, which also induces a competitive approach to prune graphs for lottery tickets at initialization.
@@ Appendix A Proofs
Now we consider again the four graphs from Figure 1. In Figure 4 we specify which nodes have which configuration from the set {A, B, C, D, E} as named in Table 8. We arbitrarily choose orange nodes to be the negative class. After one round of mean aggregation on each of them, we can estimate the amount of class information remaining on the two classes by averaging the corresponding node representations of each node per class —that is, we average the four expected means for purple nodes and the four expected means for orange nodes. We calculate these values on Table 9. As we intended to prove, they tend towards non-informative zero for both classes as the number of edges increases, and they follow the same order as the smoothing rate curves plotted in Figure 2(a). Proposition 2.3 is proved because values tend to zero —so both classes’ averages get closer together— from $\mathcal{G}^{-}$ to $\mathcal{G}$, and from $\mathcal{G}$ to both $\mathcal{G}^{+}$ and $\widetilde{\mathcal{G}^{+}}$. Proposition 2.4 is proved because the values from $\mathcal{G}^{+}$ are more informative/further apart than the values from $\widetilde{\mathcal{G}^{+}}$. ∎
@@ Appendix B How Other Label Configurations Affect the Rings’ Smoothing Rates
As a first modification (Figure 5(c)), we rotate the labels so that edge $\{0,3\}$ is now intra-class; this makes edge $\{4,7\}$ from $\widetilde{\mathcal{G}^{+}}$ intra-class, too. It is reflected on its smoothing rate plot in two main ways. First, the distance between graph $\mathcal{G}^{-}$ and $\mathcal{G}$ is not as wide, because the extra intra-edge in $\mathcal{G}$ does not cause as much smoothing as the inter-class edge from the original configuration does. Second, graph $\widetilde{\mathcal{G}^{+}}$ is now the least smoothed. This might be because the two edges aid in isolating the flow of information between the two, very distinct classes; note that this graph also has the smallest spectral gap, so the configuration of labels and the graph structure work towards the same goal.
As a second modification (Figure 5(e)), we alternate classes two by two nodes at a time. This makes $\{0,3\}$ and $\{4,7\}$ intra-class again, so it is directly comparable to the previous disposition. However, the edges in $\widetilde{\mathcal{G}^{+}}$ are not dividing the two classes so distinctively. This makes its smoothing occur more quickly than before, now en par with the base graph. We consider this phenomenon to be directly related to its lower spectral gap. Another relevant aspect of this graph is that, still, the pruned graph $\mathcal{G}^{-}$ smooths less than $\mathcal{G}$, even when the pruned edge is intra-class, and even if the spectral gap has increased; it is another instance of both mitigating over-smoothing and over-squashing.
@@ Appendix D Spectral Pruning Can Slow Down the Rate of Smoothing
In section §2.2 we have demonstrated the possibility of addressing both over-squashing and over-smoothing via spectral gap based pruning in a simple toy graph setting. Below we present the results on real-world graphs, where spectral pruning can help slow down the rate of smoothing. We adopt the same Linear GNN setup (Keriven, 2022).
In Figure 6, we present two homophilic datasets (Cora and Citeseer) and two heterophilic graphs (Texas and Chameleon). For each of these experiments we add edges using FoSR (Karhadkar et al., 2023) and ProxyAdd and delete edges using our proposed ProxyDelete method. FoSR, which optimizes the spectral gap by adding edges, aids to mitigate over-squashing but inevitably leads to accelerating the smoothing rate. Conversely, if we delete edges using our ProxyDelete method, the rate of smoothing is slowed down. It is also evident that our pruning method is more effective on heterophilous graph settings. This is likely due to the deletion of edges between nodes with different labels, thus preventing detrimental smoothing. In a recent work by Azabou et al. (2023), the authors also show similar experiments by introducing additional nodes to slow down the rate of message passing and thus slowing down the rate of smoothing. We achieve a similar effect just by pruning edges instead of introducing additional nodes.

### v2

@@ 3 Theoretical insights into spectral rewiring
Note that $\widetilde{\mathcal{G}^{+}}$ decreases the spectral gap, while $\mathcal{G}^{+}$ increases it relative to $\mathcal{G}$. However, the label configuration of $\widetilde{\mathcal{G}^{+}}$ seems more advantageous because, for the changed nodes, the number of neighbors of the same class label remains in the majority in contrast to $\mathcal{G}^{+}$. Still, the spectral gap increase seems to aid the learning task compared to the spectral gap decrease.
@@ 4 Braess-inspired graph rewiring
We introduce two algorithmic approaches to perform spectral rewiring. Our main proposal is computationally more efficient and more effective in spectral gap approximation than baselines, as we also showcase in Table 14. The other approach based on Eldan’s Lemma is also analyzed, as it provides theoretical guarantees for edge deletions. However, it does not scale well to larger graphs.
@@ 4.1 Graph rewiring with Proxy spectral gap updates
Time Complexity of ProxyAdd. The run time analysis consists of the same elements as the edge deletion algorithm. The key distinction is that the ranking is conducted on the complement of the graph’s edges, $\bar{\mathcal{E}}$. Since the set of missing edges is usually larger than the existing edges in real world settings, to save computational overhead, it is possible to only sample a constant amount of edges. See Section F for empirical runtimes.
@@ 4.2 Graph rewiring with Eldan’s criterion
Lemma 3.1 states a sufficient condition for the Braess paradox. It naturally defines a scoring function of edges to rank them according to their potential to maximize the spectral gap based on the function $g$. However, the computation of this ranking is significantly more expensive than other considered algorithms, as each scoring operation needs access to the leading eigenvector of the perturbed graph with an added or deleted edge. In case of edge deletions, we also need to approximate the spectral gap similar to our Proxy algorithms. As the involved projection $\mathcal{P}_{f}$ is a dot product of eigenvectors, it requires $\mathcal{O}(|\mathcal{V}|)$ operations. Even though this algorithm does not scale well to large graphs without focusing on a small random subset of candidate edges, we still consider it as baseline, as it defines a more conservative criterion to assess when we should stop deleting edges. The precise algorithms are stated in Appendix C.
@@ 5 Experiments
5.1 Long Range Graph Benchmark
The Long Range Graph Benchmark (LRGB) was introduced by Dwivedi et al. (2023) specifically to create a test bed for over-squashing. We compare our proposed ProxyAdd and ProxyDelete methods with DRew (Gutteridge et al., 2023), a recently proposed strong baseline for addressing over-squashing using a GCN as our backbone architecture in Table 1. We adopt the experimental setting of Tönshoff et al. (2023), we adopt DRew baseline results from the original paper. We evaluate on the following datasets and tasks: 1) PascalVOC-SP - Semantic image segmentation as a node classification task operating on superpixel graphs. 2) Peptides-func - Peptides modeled as molecular graphs. The task is graph classification. 3) Peptides-struct - Peptides modeled as molecular graphs. The task is to predict various molecular properties, hence a graph regression task.
The top performance is highlighted in bold. Evidently, our proposed rewiring methods outperform DRew (Gutteridge et al., 2023) and FoSR (Karhadkar et al., 2023) on PascalVOC and Peptides-struct, and achieves comparable performance on Peptides-func.
In addition, Table 10 in the appendix compares different rewiring strategies for node classification on other commonly used datasets and graph classification (§E.2) for adding edges, since FoSR (Karhadkar et al., 2023) was primarily tested on this task.
@@ 5.1 Long Range Graph Benchmark
Platonov et al. (2023) point out that most progress on heterophilic datasets is unreliable since many of the used datasets have drawbacks, including duplicate nodes in Chameleon and Squirrel datasets, which lead to train-test data leakage. The sizes of the small graph sizes also lead to high variance in the obtained accuracies. Consequently, we also test our proposed algorithms on 3/5 of their newly introduced larger datasets and use GCN (Kipf & Welling, 2017) and GAT (Veličković et al., 2018) as our backbone architectures. As a higher depth potentially increases over-smoothing, we also analyze how our methods fares with varied number of layers. To that end, we adopt the code base and experimental setup of Platonov et al. (2023); the datasets are divided into 50/25/25 split for train/test/validation respectively. The test accuracy is reported as an average over 10 runs. To facilitate training deeper models, skip connections and layer normalization are employed. We compare FoSR (Karhadkar et al., 2023) and our proposals based on the Eldan criterion as well as ProxyAdd and ProxyDelete in Tables 2,4,4. The top performance is highlighted in bold. Evidently, for increasing depth, even though the GNN performance should degrade because of over-smoothing, we achieve a significant boost in accuracy compared to baselines, which we attribute to the fact that our methods delete inter-class edges —thus slowing down detrimental smoothing.
@@
Table 3: Node classification on Amazon-Ratings.
@@
Table 4: Node classification on Minesweeper.
@@
In Sections §3.2 and §1, we have shown that graph pruning can improve generalization, mitigate over-squashing and also help slow down the rate of smoothing. Can we also use our insights to find lottery tickets (Frankle & Carbin, 2019)?
Method Cora Citeseer Pubmed
Metrics GS WS Acc GS WS Acc GS WS Acc
@@
To what degree is graph pruning feature data dependent? The first extension of the Lottery Ticket Hypothesis to GNNs, called Unified Graph Sparsification (UGS) (Chen et al., 2021), prunes connections in the adjacency matrix and model weights that are deemed less important for a prediction task. Note that UGS relies on information that is obtained in computationally intensive prune-train cycles that take into account the data and the associated masks. In the context of GNNs, the input graph plays a central role in determining a model’s performance at a downstream task. Naively pruning the adjacency matrix without characterizing what constitutes important edges is a pitfall we would want to avoid (Hui et al., 2023), yet resorting to expensive train-prune-rewind cycles to identify importance is also undesirable. This brings forth the questions: To what extent does the pruning criterion need to depend on the data? Is it possible to formulate a data/feature agnostic pruning criterion that optimizes a more general underlying principle to find lottery tickets? Morcos et al. (2019) and Chen et al. (2020) show, in the context of computer vision and natural language processing respectively, that lottery tickets can have universal properties that can even provably (Burkholz et al., 2022) transfer to related tasks.
Lottery tickets that rely on the spectral gap. However, even specialized structures need to maintain and promote information flow through their connections. This fact has inspired works like Pal et al. (2022); Hoang et al. (2023) to analyze how well lottery ticket pruning algorithms maintain the Ramanujan graph property of bipartite graphs, which is intrinsically related to the Cheeger constant and thus the spectral gap. They have further shown that rejecting pruning steps that would destroy a proxy of this property can positively impact the training process.
@@
In the context of GNNs, we show that we can base the graph pruning decision even entirely on the spectral gap, but rely on a computationally cheaper approach to obtain a proxy. By replacing the magnitude pruning criterion for the graph with the Eldan criterion and ProxyDelete to prune edges, in principle, we can avoid the need for additional data features and labels. This has the advantage that we could also prune the graph at initialization and thus benefit from the computational savings from the start. We use our proposed methods to prune the graph at initialization to the requisite sparsity level and then feed it to the GNN where the weights are pruned in an iterative manner. Our results are presented in Table 18, where we compare IMP based UGS (Chen et al., 2021) with our methods for different graph and weight sparsity levels. Note that, even though our method does not take any feature information into account and prunes purely based on the graph structure, our results are comparable. For datasets like Pubmed, we even slightly outperform the baseline. Table 5 shows results for jointly pruning the graph and parameter weights, which leads to better results due to potential positive effects of overparameterization on training (Gadhikar & Burkholz, 2024).
@@
Stopping criterion. The advantage of using spectral gap based pruning (especially the Eldan criterion) is patent: It helps identify problematic edges that cause information bottlenecks and provides a framework to prune those edges. Unlike UGS, our proposed framework also has the advantage that we can couple the overall pruning scheme with a stopping criterion that follows naturally from our setup. We stop pruning the input graph when no available edges satisfy our criterion anymore.
@@ 6 Conclusion
Our work connects two seemingly distinct branches of the literature on GNNs: rewiring graphs to mitigate over-squashing and pruning graphs for lottery tickets to save computational resources. Contributing to the first branch, we highlight that, contrary to the standard rewiring practice, not only adding but also pruning edges can increase the spectral gap of a graph exploiting the Braess paradox. By providing a minimal example, we prove that this way it is possible to address over-squashing and over-smoothing simultaneously. Experiments on large-scale heterophilic graphs confirm the practical utility of this insight. Contributing to the second branch, these results explain how pruning graphs moderately can improve the generalization performance of GNNs, in particular for heterophilic learning tasks. To utilize these insights, we have proposed a computationally efficient graph rewiring framework, which also induces a competitive approach to prune graphs for lottery tickets at initialization.
@@ Appendix A Proofs
Now we consider again the four graphs from Figure 1. In Figure 4 we specify which nodes have which configuration from the set {A, B, C, D, E} as named in Table 7. We arbitrarily choose orange nodes to be the negative class. After one round of mean aggregation on each of them, we can estimate the amount of class information remaining on the two classes by averaging the corresponding node representations of each node per class —that is, we average the four expected means for purple nodes and the four expected means for orange nodes. We calculate these values in Table 8. As we intended to prove, they tend towards non-informative zero for both classes as the number of edges increases, and they follow the same order as the smoothing rate curves plotted in Figure 2(a). Proposition 3.3 is proved because values tend to zero —so both classes’ averages get closer together— from $\mathcal{G}^{-}$ to $\mathcal{G}$, and from $\mathcal{G}$ to both $\mathcal{G}^{+}$ and $\widetilde{\mathcal{G}^{+}}$. Proposition 3.4 is proved because the values from $\mathcal{G}^{+}$ are more informative/further apart than the values from $\widetilde{\mathcal{G}^{+}}$. ∎
@@ Appendix B How other label configurations affect the rings’ smoothing rates
As a first modification (Figure 5(c)), we rotate the labels so that edge $\{0,3\}$ is now intra-class; this makes edge $\{4,7\}$ from $\widetilde{\mathcal{G}^{+}}$ intra-class, too. It is reflected in its smoothing rate plot in two main ways. First, the distance between graph $\mathcal{G}^{-}$ and $\mathcal{G}$ is not as wide, because the extra intra-edge in $\mathcal{G}$ does not cause as much smoothing as the inter-class edge from the original configuration does. Second, graph $\widetilde{\mathcal{G}^{+}}$ is now the least smoothed. This might be because the two edges aid in isolating the flow of information between the two, very distinct classes; note that this graph also has the smallest spectral gap, so the configuration of labels and the graph structure work towards the same goal.
As a second modification (Figure 5(e)), we alternate classes two by two nodes at a time. This makes $\{0,3\}$ and $\{4,7\}$ intra-class again, so it is directly comparable to the previous disposition. However, the edges in $\widetilde{\mathcal{G}^{+}}$ are not dividing the two classes so distinctively. This makes its smoothing occur more quickly than before, now on par with the base graph. We consider this phenomenon to be directly related to its lower spectral gap. Another relevant aspect of this graph is that, still, the pruned graph $\mathcal{G}^{-}$ smooths less than $\mathcal{G}$, even when the pruned edge is intra-class, and even if the spectral gap has increased; it is another instance of both mitigating over-smoothing and over-squashing.
@@ Appendix D Spectral pruning can slow down the rate of smoothing
In section §3.2 we have demonstrated the possibility of addressing both over-squashing and over-smoothing via spectral gap based pruning in a simple toy graph setting. Below we present the results on real-world graphs, where spectral pruning can help slow down the rate of smoothing. We adopt the same Linear GNN setup (Keriven, 2022). In Figure 6, we present two homophilic datasets (Cora and Citeseer) and two heterophilic graphs (Texas and Chameleon). For each of these experiments we add edges using FoSR (Karhadkar et al., 2023) and ProxyAdd and delete edges using our proposed ProxyDelete method. FoSR, which optimizes the spectral gap by adding edges, aids in mitigating over-squashing but inevitably leads to accelerating the smoothing rate. Conversely, if we delete edges using our ProxyDelete method, the rate of smoothing is slowed down. It is also evident that our pruning method is more effective in heterophilous graph settings. This is likely due to the deletion of edges between nodes with different labels, thus preventing detrimental smoothing. We substantiate this by measuring the distance between nodes that have different labels, which should stay distinguishable. That is, our method deletes edges between nodes of different labels thus preventing unnecessary aggregation. We report the cosine distance for heterophilic graphs in Table 9 before training, after training on the original graph, and after training on the pruned graph. From the table it is clear that pruning edges increases the distance between nodes of different labels. Another popular metric in the literature to measure over-smoothing is Dirichlet energy, which can only measure the degree of smoothing, but not whether it is helpful for a learning task. To keep up with the trend, we plot the Dirichlet energy vs. Layers (Roth & Liebig, 2023) in Figure 7 on Cora and Texas. It is clear from the figure that our method slows down the decay of Dirichlet energy. Note that, since our method works purely on the graph topology, it cannot improve the Dirichlet energy like specialised methods (Zhou et al., 2021; Roth & Liebig, 2023; Rusch et al., 2023b).
In a recent work by Azabou et al. (2023), the authors also show similar experiments by introducing additional nodes to slow down the rate of message passing and thus slowing down the rate of smoothing. We achieve a similar effect just by pruning edges instead of introducing additional nodes.
@@
Table 9: Cosine distance between nodes of different labels before and after deleting edges using ProxyDelete.
Dataset
Before
Training
After
Training
After
Training
Cornell 0.72 0.87 0.83
Wisconsin 0.72 0.77 0.86
Texas 0.68 0.62 0.80
Chameleon 0.99 0.91 0.96
Squirrel 0.98 0.82 0.89
Actor 0.83 0.95 0.99
(a) Cora dataset with 50 edges added (FoSR) and 50 deleted (ProxyDelete).
(b) Texas dataset with 50 edges added (FoSR) and 100 deleted (ProxyDelete) .
Figure 7: We measure the Dirichlet energy and plot it for increasing depth on a homophilic dataset, Cora and a heterophilic dataset, Texas. For increasing depth, we can see Proxydelmax slows the decay of Dirichlet energy.